A 1200 kg car downhill speeds from rest to 2.4m/s² (acceleration) uphill in 14 seconds to reach the top of the hill. The height of the hill is 48m. If the car is 31% efficient, how much energy did the car use?
1
Expert's answer
2019-11-21T10:06:42-0500
"E_p=mgh"

"E=\frac{E_p}{\eta}"

"E=\frac{(1200)(9.8)(48)}{0.31}=1.82\cdot 10^6\ J"